20190107 Measure new DX Flat with FLAO Closed loop

Startup DX system

16:42 SetFlat AO
  • Act 341 position 84.22 um
  • 72 modes
    • Act 342 0.304N
  • Using default.sav -> ../20180912_021523/flattening_result.sav
  • Tried flat/Applicatable/20190105_400_mode_3.sav measured by Greg. 74 modes

17:10 Power on WUnit

17:55 Rest and Set
  • 326 modes applied.
  • default flat:
    pupils_flat_default.jpg

19:00 CLosed loop, bin 1, 12 modes
  • TT gain, OK
  • HO1 gain, looks like pupils are rotating
  • Checked, re-rotator is running.
  • Stop and start DX WFS and AdSec.
  • Try again

20:06 Looks like the wrong pupils on the pyramid were used
  • did not complain when I loaded the b1m12 recon.
  • change_binning.py 1
    • 1 = 171742
  • 12 modes OK
  • 400 modes @ 990 HZ OK
  • Flat 20190107_400_1
  • Open Loop
  • Set Flat with 20190107_400_1
    • 343 modes
    • Act 454 0.2 N
    • Act 450 -0.282 (next to 454 which is lower outer edge between two acts with no position)

20:18 Try flat with 400 modes and 200 Hz with 6.0 Lam/D
  • image

20:22 500Hz, 6 lam/D
  • image

20:30 990 Hz, 3 lam/D
  • 990 Hz Closed Loop Pupils:
    pupils_990Hz_3LoD.jpg
  • 990 Hz Closed Loop Mirror GUI:
    mirror_990Hz_3LoD.jpg
  • flat with 20190107_400_1, 327 modes applied:
    pupils_flat_20190107_400_1.jpg

20:53 Final Setup
  • Bayside Stages -72, -49.3, 55.6
  • Hexapod Position:
    hex_final.jpg

There is still a small bite out of the the pupils at 5:30-6:00. This is where we are missing many (7 of 10) actuators along the bottom of the DX AdSec. These 7 actuators were repaired last summer but we do not yet have reconstructors for FLAO, ARGOS nor LBTI for the new actuator configuration. We are working on FLAO reconstructors now (last 3 days and again in the next couple days)

Pat moved the telescope to 50 Elevation so I can test the new flat.
  • 90 El -0.28, 0.19 N max/min
  • 70 El -0.30, 0.19 N max/min
  • 50 El -0.37, 0.39 N Max/min (Act 454 at the lower, outer diameter and Act 380 flutuating/Acts 217, 216, 104 at the inner diameter)

21:19 Set flat at 50 El
  • -0.35, 0.12 N max/min (Act 380 /Act 573 with bump around
  • 138 modes applied.
  • Flat applied at 50 Elevation:
    flat_50_El.jpg

  • Flat
  • Flat 570 cause 0 modes applied
  • Flat 142 modes, Act 569
  • Flat 573 cause 0 modes applied
  • Flat 138 modes, Act 582

We found that only ~140 modes are applied at 50 Elevation, with high (~-0.3 N) forces at the top edge of the central membrane. Guido and I are concerned about these high forces (black in the center) and will look in the telemetry to see if this is a new feature (it did exist in late December) or has been there since last February when TS3 was re-installed after re-coating.

-- DougMiller - 07 Jan 2019

I Attachment Action Size Date Who Comment
flat_50_El.jpgjpg flat_50_El.jpg manage 222 K 07 Jan 2019 - 21:29 DougMiller Flat applied at 50 Elevation
hex_final.jpgjpg hex_final.jpg manage 29 K 07 Jan 2019 - 20:52 DougMiller Hexapod Position
mirror_990Hz_3LoD.jpgjpg mirror_990Hz_3LoD.jpg manage 340 K 07 Jan 2019 - 20:43 DougMiller 990 Hz Closed Loop Mirror GUI
pupils_990Hz_3LoD.jpgjpg pupils_990Hz_3LoD.jpg manage 62 K 07 Jan 2019 - 20:42 DougMiller 990 Hz Closed Loop Pupils
pupils_flat_20190107_400_1.jpgjpg pupils_flat_20190107_400_1.jpg manage 60 K 07 Jan 2019 - 20:44 DougMiller flat with 20190107_400_1, 327 modes applied
pupils_flat_default.jpgjpg pupils_flat_default.jpg manage 58 K 07 Jan 2019 - 20:41 DougMiller default flat
Topic revision: r2 - 07 Jan 2019, DougMiller
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback